solve each for x :
( Can you please provide steps and explanations. Thank you)
1.) 1/4=8^(x+3)
2.) 27^(x+1)=9^(2x-4)

solve for x in each:
( Can you pleas provide steps and explanation. I would really appreciate it. It would help me to better understand the process in obtaining the answers.)
1.)log base x of 8 = 3/2
2.)log base a of x = 3
3.)ln e^(x)= 4

Answer by lwsshak3(11628)

solve each for x :
( Can you please provide steps and explanations. Thank you)
1.) 1/4=8^(x+3)
2.) 27^(x+1)=9^(2x-4)
solve for x in each:
( Can you pleas provide steps and explanation. I would really appreciate it. It would help me to better understand the process in obtaining the answers.)
1.)log base x of 8 = 3/2
2.)log base a of x = 3
3.)ln e^(x)= 4
**
1.) 1/4=8^(x+3)
make the base same on both sides
1/2^2=2^3^(x+3)
2^-2=2^3(x+3)
equate exponents with same base(2)
-2=3(x+3)=3x+9
3x=-11
x=-11/3
..
2.) 27^(x+1)=9^(2x-4)
3^3^(x+1)=3^2^(2x-4)
3^3(x+1)=3^2(2x-4)
equate exponents
3(x+1)=2(2x-4)
3x+3=4x-8
x=11
..
1.)log base x of 8 = 3/2
convert to exponential form: (base(x) raised to log of number(3/2)=number(8)
x^(3/2)=8
raise both sides by 2/3
x^(3/2)^(2/3)=8^(2/3)
x=8^2/3=4
..
2.)log base a of x = 3
convert to exponential form
a^3=x
..
3.)ln e^(x)= 4
base(e) raised to a power(x)=x
x=4